Even Integers. If the sum of two consecutive integers is 86, what is the two integers.
Answer by Earlsdon(6294)

Let x = the first even integer, then x+2 = the next consecutive even integer.
The sum: x + x+2 = 86 Collect like-terms.
2x + 2 = 86 Subtract 2 from both sides.
2x = 84 Divide both sides by 2.
x = 42.
The two consecutive even integers are: 42 and 44.
Check: 42 + 44 = 86